# Covariant effective spacetimes of spherically symmetric electro-vacuum with a cosmological constant

### v1

@@ III.3 Solutions to effective dynamics from $M_{\rm eff,I}$
As a consequence, $\mathcal{X}(\mu_{\rm I})$ and $\mathcal{Y}(\mu_{\rm I})$ are two constants. Without loss of generality, we set $\mathcal{X}(\mu_{\rm I})=\mathcal{Y}(\mu_{\rm I})=1$, since any other non-vanishing values of $\mathcal{X}(\mu_{\rm I})$ and $\mathcal{Y}(\mu_{\rm I})$ would simply correspond to redefinitions of $\Lambda$ and $Q$.
@@
$-x^{2}\left[\frac{{\rm i}\zeta e^{\frac{{\rm i}2\zeta K_{2}}{x}}}{(E^{2})^{2}}-\frac{\sin\left(\frac{2\zeta K_{2}}{x}\right)}{2\zeta}\right]=0.$ (117)
Combing Eq. (117) with Eq. (III.3), one can eliminate $K_{2}$ to get
$E^{2}$ =±$=\pm\frac{x}{\sqrt{\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\left[1+\frac{\zeta^{2}}{x^{2}}\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\right]}}.$ (118)
@@
$=f_{c}(x)\left[1+\frac{\zeta^{2}}{x^{2}}f_{c}(x)\right].$ (121)
The solution (120) reduces to the one obtained in [50] as $Q\rightarrow 0$, and to the vacuum one in [41, 42] as $Q=\Lambda\rightarrow 0$. We can also choose the PG gauge (73) to get the metric. In this gauge, the lapse function becomes to $N=1$, according to Eq. (114). Combining Eq. (107a) with Eq. (113) yields
@@
Hence, these two vector potentials obtained in these two different gauges are the same up to the gauge term ${\rm d}Y_{\rm eff,I}(x)$, reflecting the covariance of EM field.
@@ III.4 Solutions to effective dynamics from $M_{\rm eff,II}$
To determine resulting solution, one needs to make a certain gauge fixing. In the Schwarzschild gauge (67), Eq. (107a) leads to
@@
Combing Eqs. (129) and (III.4), we have
$E^{2}$ =±$=\pm\frac{x}{\sqrt{1-\frac{2GM_{\rm eff,II}}{x}}}$
±$\equiv\pm\frac{x}{\sqrt{f_{\rm eff,II}(x)}}.$ (130)
@@
${\rm d}s^{2}$ $=-e^{g_{\rm eff,II}(x)}f_{\rm eff,II}(x){\rm d}t^{2}+\frac{{\rm d}x^{2}}{\mu_{\rm II}(x)f_{\rm eff,II}(x)}+x^{2}{\rm d}\Omega^{2}.$ (132)
Similarly, in the PG gauge (73), combing Eqs. (107a) and (III.4), we have
$N^{x}$ =±$=\pm N\sqrt{\mu_{\rm II}\frac{2GM_{\rm eff,II}}{x}}$
=±$=\pm N\sqrt{\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}.$ (133)
@@
On the other hand, Eq. (109) can be written as.
@@
Hence, the solution to Eq. (137) is can be expressed as
@@
indicating the covariance of EM field.
In the following, we will calculate the values $M_{\rm eff,II}$, $g_{\rm eff,II}(x)$ and $\Psi_{\rm eff,II}(x)$ for different choices of $\mathcal{X}(\mu_{\rm II})$ and $\mathcal{Y}(\mu_{\rm II})$.
@@ III.4.2 Case of $\mathcal{X}(\mu_{\rm II})=1$ and $\mathcal{Y}(\mu_{\rm II})=\mu_{\rm II}$
Now we calculate $\Psi_{\rm eff,II}$ from Eq. (138), and obtain
@@ III.5 Solutions to effective dynamics from $M_{\rm eff,III}$
$\mu_{\rm III}$ $=1-\frac{4\zeta^{4}(GM_{\rm eff,III})^{2}}{(s_{1})^{3}}$
$=1-\frac{4\zeta^{4}(GM_{\rm eff,III})^{2}}{x^{6}}.$ (158)
@@
Combing Eqs. (159) and (157), we obtain
$E^{2}$ =±$=\pm\frac{x}{\sqrt{1-\frac{x^{2}}{\zeta^{2}}\left[n\pi+(-1)^{n}\arcsin\left(\frac{2\zeta^{2}GM_{\rm eff,III}}{x^{3}}\right)\right]}}$
±$\equiv\pm\frac{x}{\sqrt{f_{\rm eff,III}(x)}}.$ (160)
@@
$N^{2}$ $=\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{Gx^{2}}{\rm d}x\right]f_{\rm eff,III}(x)$
$\equiv e^{g_{\rm eff,III}(x)}f_{\rm eff,III}(x).$ (161)
@@
${\rm d}s^{2}$ $=-e^{g_{\rm eff,III}(x)}f_{\rm eff,III}(x){\rm d}t^{2}+\frac{{\rm d}x^{2}}{\mu_{\rm III}(x)f_{\rm eff,III}(x)}+x^{2}{\rm d}\Omega^{2}.$ (162)
Similarly, in the PG gauge (73), combing Eqs. (107a) and (157), we obtain
$N^{x}$ =±$=\pm N\sqrt{\mu_{\rm III}(x)\left[1-f_{\rm eff,III}(x)\right]}.$ (163)
@@
${\rm d}s^{2}=$ $-e^{g_{\rm eff,III}(x)}{\rm d}\tau^{2}$
$+\frac{1}{\mu_{\rm III}(x)}\left({\rm d}x\pm\sqrt{e^{g_{\rm eff,III}(x)}\mu_{\rm III}(x)\left[1-f_{\rm eff,III}(x)\right]}{\rm d}\tau\right)^{2}$
@@
The remaining calculations and the results are the same as those in Sec. III.4, just by replacing subscript II by III.

### v3

@@ III.3 Solutions to effective dynamics from $M_{\rm eff,I}$
As a consequence, $\mathcal{X}(\mu_{\rm I})$ and $\mathcal{Y}(\mu_{\rm I})$ are two constants. Without loss of generality, we set $\mathcal{X}(\mu_{\rm I})=\mathcal{Y}(\mu_{\rm I})=1$, since any other nonvanishing values of $\mathcal{X}(\mu_{\rm I})$ and $\mathcal{Y}(\mu_{\rm I})$ would simply correspond to redefinitions of $\Lambda$ and $Q$.
@@
$x^{2}\left[\frac{{\rm i}\zeta e^{\frac{{\rm i}2\zeta K_{2}}{x}}}{(E^{2})^{2}}-\frac{\sin\left(\frac{2\zeta K_{2}}{x}\right)}{2\zeta}\right]=0.$ (104)
Combining Eq. (104) with Eq. (III.3), one can eliminate $K_{2}$ to get
$E^{2}$ $=\pm\frac{x}{\sqrt{\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\left[1+\frac{\zeta^{2}}{x^{2}}\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\right]}}.$ (105)
@@
$=f_{\rm cl}(x)\left[1+\frac{\zeta^{2}}{x^{2}}f_{\rm cl}(x)\right].$ (108)
The solution (107) reduces to the one obtained in [56] as $Q\rightarrow 0$ and to the vacuum one in [42, 43] as $Q=\Lambda\rightarrow 0$. We can also choose the PG gauge (60) to get the metric. In this gauge, the lapse function becomes $N=1$, according to Eq. (101). Combining Eq. (94a) with Eq. (100) yields
@@
Hence, these two vector potentials obtained in these two different gauges are the same up to the gauge term ${\rm d}Y_{\rm eff,I}(x)$, reflecting the covariance of the EM field.
III.3.1 Properties of spacetime ${\rm d}s^{2}_{\rm eff,I}$ with $\Lambda=0$
To gain a preliminary insight into the structure of the spacetime, we assume $\Lambda=0$ for simplicity. In addition, we will focus on the physically interesting case where $GM\gg\zeta$, $Q\ll\sqrt{G}M$, corresponding to a black hole with a large mass and a small charge. In this case, $f_{\rm cl}$ will have two distinct roots $x_{-}<x_{+}$, respectively, corresponding to the inner horizon $x_{-}$ and the outer horizon $x_{+}$ in the classical theory.
Now let us focus on the horizons of the effective spacetime ${\rm d}s^{2}_{\rm eff,I}$. According to Eq. (108), $f_{\rm eff,I}$ can be factorized such that $f_{\rm cl}$ appears as one of its factors. It then follows that $x_{\pm}$ remain roots of $f_{\rm eff,I}$. Namely, $x=x_{\pm}$ still correspond to horizons in the effective spacetime described by ${\rm d}s^{2}_{\rm eff,I}$. For the second factor $1+\frac{\zeta^{2}}{x^{2}}f_{\rm cl}(x)$, we note that its roots are the intersection of the graph of $y=\frac{\zeta^{2}}{x^{2}}f_{\rm cl}(x)$ with the horizontal line y=1y=-1. The qualitative behavior of $\frac{\zeta^{2}}{x^{2}}f_{\rm cl}(x)$ can be understood from the following facts:
  • (1)
    It shares the same roots $x_{\pm}$ as $f_{\rm cl}(x)$.
  • (2)
[…]
  • (3)
    The value of the function at $(x_{+}+x_{-})/2=Q^{2}/M$ is less than 1-1 in the regime under consideration.
These facts, together with the expression of $f_{\rm cl}$, allow us to determine the graph of $\frac{\zeta^{2}}{x^{2}}f_{\rm cl}$. As shown in Fig. 1, the function $1+\frac{\zeta^{2}}{x^{2}}f_{\rm cl}(x)$ has two distinct roots, denoted by $x^{\prime}_{\pm}$, satisfying $x_{-}<x^{\prime}_{-}<x^{\prime}_{+}<x_{+}$. As a consequence, the effective spacetime ${\rm d}s^{2}_{\rm eff,I}$ possesses two additional horizons located at $x=x^{\prime}_{\pm}$, which lie between the classical inner and outer horizons. A further analysis of the spacetime structure involves addressing the nontrivial problem of constructing a Penrose diagram for a spacetime with more than one inner horizon [57]. This issue will be explored in future work.
Figure 1: The graph of $f_{\rm cl}$ (red solid line), $\frac{\zeta^{2}}{x^{2}}f_{\rm cl}$ (black solid line), and the horizonal line y=1y=-1 (black dashed line). According to the facts (1) and (3) given in Sec. III.3.1, the graphs pass by the points $(x_{\pm},0)$ (black dot) and $(Q^{2}/M,y_{o})$ (red dot) with $y_{o}<-1$. In addition, no unexpected features such as oscillations appear in $\frac{\zeta^{2}}{x^{2}}f_{\rm cl}$, since it is simply rescaled from $f_{\rm cl}$, which varies smoothly with a single minimum.
@@ III.4 Solutions to effective dynamics from $M_{\rm eff,II}$
To determine the resulting solution, one needs to make a certain gauge fixing. In the Schwarzschild gauge (54), Eq. (94a) leads to
@@
Combining Eqs. (116) and (III.4), we have
$E^{2}$ $=\pm\frac{x}{\sqrt{1-\frac{2GM_{\rm eff,II}}{x}}}.$ (117)
@@
${\rm d}s^{2}_{\rm eff,II}$ $=-e^{g_{\rm eff,II}(x)}f_{\rm eff,II}(x){\rm d}t^{2}+\frac{{\rm d}x^{2}}{\mu_{\rm II}(x)f_{\rm eff,II}(x)}+x^{2}{\rm d}\Omega^{2}.$ (119)
Similarly, in the PG gauge (60), combining Eqs. (94a) and (III.4), we have
$N^{x}$ $=\pm N\sqrt{\mu_{\rm II}\frac{2GM_{\rm eff,II}}{x}}$
$=\pm N\sqrt{\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}.$ (120)
@@
On the other hand, Eq. (96) can be written as
@@
Hence, the solution to Eq. (124) can be expressed as
@@
indicating the covariance of the EM field.
In the following, we will calculate the values $M_{\rm eff,II}$, $g_{\rm eff,II}(x)$, and $\Psi_{\rm eff,II}(x)$ for different choices of $\mathcal{X}(\mu_{\rm II})$ and $\mathcal{Y}(\mu_{\rm II})$, which have exact analytical solutions.
@@ III.4.2 Case of $\mathcal{X}(\mu_{\rm II})=1$ and $\mathcal{Y}(\mu_{\rm II})=\mu_{\rm II}$
Now we calculate $\Psi_{\rm eff,II}$ from Eq. (125) and obtain
@@ III.4.3 Case of $\mathcal{X}(\mu_{\rm II})=\mu_{\rm II}$ and $\mathcal{Y}(\mu_{\rm II})=1$
III.4.4 Properties of spacetime ${\rm d}s_{\rm eff,II}^{2}$ with $\Lambda=0$
In this section, we analyze the properties of the spacetime ${\rm d}s_{\rm eff,II}^{2}$, using the case presented in Sec. III.4.1 as a concrete example. For the case discussed in Secs. III.4.2 and III.4.3, the analysis is similar but technically more involved.
For simplicity, we again assume $\Lambda=0$, $GM\gg\zeta$, and $Q\ll\sqrt{G}M$. According to Eq. (119), the metric ${\rm d}s_{\rm eff,II}^{2}$ under the Schwarzschild-like coordinates takes the form
${\rm d}s_{\rm eff,II}^{2}=-f(x){\rm d}t^{2}+g(x)^{-1}{\rm d}x^{2}+x^{2}{\rm d}\Omega^{2},$ (144)
with $f\neq g$. In the case we considered, $f$ is proportional to $f_{\rm cl}$ by a nonvanishing factor. Thus, as discussed in Sec. III.3.1, it possesses two positive roots: $x_{\pm}$. Moreover, $g(x)$ is equal to $f_{\rm eff,I}$, and consequently, it owns four positive roots: $x_{\pm}$ and $x^{\prime}_{\pm}$ with $x_{-}<x_{-}^{\prime}<x_{+}^{\prime}<x_{+}$ (see Fig. 1). Although all of the roots give coordinate singularities, only the roots of $f(x)$, namely, $x_{\pm}$, correspond to horizons. According to the discussion in Sec. III.3.1, we get $f(x^{\prime}_{+})<0$, implying that the hypersurface $x=x^{\prime}_{+}$ is spacelike. Indeed, one may regard $x=x^{\prime}_{+}$ as the limit of a series of hypersurfaces $x=c$ as $x\to x^{\prime}_{+}$. In this limit, the induced metric on $x=x^{\prime}_{+}$ is given by
${\rm d}s^{2}\big|_{x=x^{\prime}_{+}}=\lim_{c\to x^{\prime}_{+}}{\rm d}s^{2}\big|_{x=c}=-f(x^{\prime}_{+}){\rm d}t^{2}+(x^{\prime}_{+})^{2}{\rm d}\Omega^{2},$ (145)
which clearly confirms its spacelike character.
Before investigating the extension of the spacetime ${\rm d}s_{\rm eff}^{2}$, it should be stressed that the domain $x>0$ is divided into five separate regions by the four coordinate singularities $_{\pm}, x±x^{\prime}_{\pm}$, and the maximal extension of one of these regions does not necessarily include the others. In our analysis, we begin with the region $x>x_{+}$. Since $x=x_{+}$ corresponds to a coordinate singularity, extending the spacetime across this surface requires a new coordinate chart that remains regular at $x=x_{+}$. The PG-like coordinate system defined by Eq. (III.4) provides such a coordinate chart. With this coordinate, the spacetime will be extended to include the region $x^{\prime}_{+}<x<x_{+}$. To extend the space beyond $x=x^{\prime}_{+}$, we follow the method suggested in [51] for a similar context. Specifically, we introduce a new coordinate $z(x)$ defined via
$\left(\frac{{\rm d}z(x)}{{\rm d}x}\right)^{2}=f(x)g(x)^{-1}=\mu_{\rm eff,II}(x)^{-1}.$ (146)
In terms of the new coordinates $(t,z)$, the metric takes the form
${\rm d}s^{2}=-f(z){\rm d}t^{2}+f(z)^{-1}{\rm d}z^{2}+x(z)^{2}{\rm d}\Omega^{2},$ (147)
where $x(z)$ is the inverse function of $z(x)$ determined in Eq. (146). Since $x=x^{\prime}_{+}$ is a root of $g(x)$, the function $x(z)$ exhibits a turning point at $x(z)=x^{\prime}_{+}$. This implies that the extension beyond $x=x^{\prime}_{+}$ does not penetrate into the region with $x<x_{+}$, but instead reflects back. Since the classical timelike singularity lies within the region $x\in(0,x_{-})$ where $x_{-}<x^{\prime}_{+}$, the spacetime in our model avoids this singularity and is thus nonsingular. The Penrose diagram corresponds to that proposed for the second model in our previous work [42], which resembles the gluing of two Kruskal spacetimes across a black-to-white hole transition surface.
@@ III.5 Solutions to effective dynamics from $M_{\rm eff,III}$
$\mu_{\rm III}$ $=1-\frac{4\zeta^{4}(GM_{\rm eff,III})^{2}}{(s_{1})^{3}}=1-\frac{4\zeta^{4}(GM_{\rm eff,III})^{2}}{x^{6}}.$ (149)
@@
Combining Eqs. (150) and (148), we obtain
$E^{2}$ $=\pm\frac{x}{\sqrt{1-\frac{x^{2}}{\zeta^{2}}\left[n\pi+(-1)^{n}\arcsin\left(\frac{2\zeta^{2}GM_{\rm eff,III}}{x^{3}}\right)\right]}}$ (151)
@@
$N^{2}$ $=\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{Gx^{2}}{\rm d}x\right]f_{\rm eff,III}^{(n)}(x)$
$\equiv e^{g_{\rm eff,III}(x)}f_{\rm eff,III}^{(n)}(x),$ (152)
where we introduced $f_{\rm eff,III}^{(n)}(x)$ as
$f_{\rm eff,III}^{(n)}(x)=1-\frac{x^{2}}{\zeta^{2}}\left[n\pi+(-1)^{n}\arcsin\left(\frac{2\zeta^{2}GM_{\rm eff,III}}{x^{3}}\right)\right].$ (153)
@@
${\rm d}s^{2}_{\rm eff,III}$ $=-e^{g_{\rm eff,III}(x)}f_{\rm eff,III}^{(n)}(x){\rm d}t^{2}+\frac{{\rm d}x^{2}}{\mu_{\rm III}(x)f_{\rm eff,III}^{(n)}(x)}+x^{2}{\rm d}\Omega^{2}.$ (154)
Similarly, in the PG gauge (60), combining Eqs. (94a) and (148), we obtain
$N^{x}$ $=\pm N\sqrt{\mu_{\rm III}(x)\left[1-f_{\rm eff,III}(x)\right]}.$ (155)
@@
${\rm d}s^{2}_{\rm eff,III}=$ $-e^{g_{\rm eff,III}(x)}{\rm d}\tau^{2}$
$+\frac{1}{\mu_{\rm III}(x)}\left({\rm d}x\pm\sqrt{e^{g_{\rm eff,III}(x)}\mu_{\rm III}(x)\left[1-f_{\rm eff,III}^{(n)}(x)\right]}{\rm d}\tau\right)^{2}$
@@
The remaining calculations are the same as those in Sec. III.4, just by replacing subscript II by III.
III.5.1 Properties of spacetime ${\rm d}s^{2}_{\rm eff,III}$ with $\Lambda=0$
For concrete analysis, we will consider the case with $\mathcal{X}=1=\mathcal{Y}$, leading to $N=1$, and
$M_{\rm eff,III}=M-\frac{Q^{2}}{2x}+\frac{\Lambda x^{3}}{6G}.$ (158)
Let us still focus on the case where $\Lambda=0$, $\sqrt{G}Q\ll GM$, and $GM\gg\zeta^{2}$.
For the metric ${\rm d}s_{\rm eff,III}^{2}$ in Schwarzschild-like coordinates, as given in Eq. (154), it retains the same form as Eq. (144). Consequently, the procedure for extending the spacetime follows exactly the same steps as those described in Sec. III.4.4.
To investigate the maximally extended spacetime of ${\rm d}s_{\rm eff,III}^{2}$, let us start with $n=0$. Due to the presence of the arcsine function in $f_{\rm eff,III}^{(0)}$, $x$ must lie within the domain $x_{\rm min},\infty)$ where $x_{\rm min}$ is defined by
$\frac{2\zeta^{2}GM_{\rm eff,III}(x_{\rm min})}{x_{\rm min}^{3}}=1.$ (159)
Within this domain we have
$\mu(x)>0,\ \forall x>x_{\rm min},\quad\text{and}\quad\mu(x_{\rm min})=0.$ (160)
In the regime with $\Lambda=0$, $\sqrt{G}Q\ll GM$, and $GM\gg\zeta^{2}$, $f_{\rm eff,III}^{(0)}$ possesses two roots, denoted by $x_{+}>x_{-}$. Due to Eq. (160), the PG coordinates given in (III.5) are well-defined across $x=x_{\pm}$. Thus, the extension of the region $x>x_{+}$ can be achieved by using the PG coordinates, and the resulting extension will cover the entire domain $x>x_{\rm min}$. To extend the spacetime beyond $x=x_{\rm min}$, we again introduce the coordinate $z(x)$ such that
$\left(\frac{{\rm d}z}{{\rm d}x}\right)^{2}=\mu_{\rm III}^{-1}.$ (161)
Then, as in Sec. III.4.4, $x(z)$ exhibit a turning point at $x(z)=x_{\rm min}$. However, unlike the case discussed in Sec. III.4.4, a subtle issue arises here due to the presence of the arcsine function in $f_{\rm eff,III}^{(n)}$: after the turning point, the line element should transition to the branch of the arcsine function corresponding to $n=1$. This transition is essential to ensure the smoothness of the metric across the turning point. Indeed, by choosing the $n=1$ branch after the turning point, the metric, in the $(t,z)$ coordinates, reads
${\rm d}s^{2}_{\rm eff,III}=\left\{\begin{array}[]{cc}-f_{\rm eff,III}^{(0)}(z){\rm d}t^{2}+\frac{{\rm d}x^{2}}{f_{\rm eff,III}^{(0)}(z)}+x(z)^{2}{\rm d}\Omega^{2},&z\geq z_{o}\\ -f_{\rm eff,III}^{(1)}(z){\rm d}t^{2}+\frac{{\rm d}x^{2}}{f_{\rm eff,III}^{(1)}(z)}+x(z)^{2}{\rm d}\Omega^{2},&z\leq z_{o}\end{array}\right$. (162)
where $z_{o}$ is the value of $z$ at the turning point and $z>z_{o}$ corresponds to the region before the turning point. Then, the smoothness of the metric is ensured due to
$\lim_{z\to z_{o}^{+}}\frac{{\rm d}^{n}}{{\rm d}z^{n}}f_{\rm eff,III}^{(0)}=\lim_{z\to z_{o}^{-}}\frac{{\rm d}^{n}}{{\rm d}z^{n}}f_{\rm eff,III}^{(1)}.$ (163)
In addition, following a similar line of reasoning, we find that if the arcsine function is constrained to remain in the $n=0$ branch after the turning point, the resulting metric becomes nondifferentiable due to
$\lim_{z\to z_{o}^{+}}\frac{{\rm d}}{{\rm d}z}f_{\rm eff,III}^{(0)}=-\lim_{z\to z_{o}^{-}}\frac{{\rm d}}{{\rm d}z}f_{\rm eff,III}^{(0)}.$ (164)
This demonstrates the necessity of transitioning to the $n=1$ branch to ensure the smoothness of the metric. The space is complete after the one corresponding to $n=1$ is included. The resulting Penrose diagram coincides with the one presented in our previous work [43], resembling that of the Kruskal spacetime but with the central singularity replaced by an asymptotically de Sitter region. This de Sitter region arises from the presence of the $n\pi$ term, which is nonvanishing for $n=1$, in Eq. (153).